# Rethinking the Value of Agent-Generated Tests for LLM-Based Software Engineering Agents

### v1

@@ 4. RQ2: What Feedback Signals Do Agent-Written Tests Provide, and What Types of Assertions Do They Use?
RQ2 conditions on tasks that write at least one test artifact and reports descriptive summaries of two aspects of test feedback:
  • Signal counts (RQ2.1): how many feedback statements appear in agent-written tests, split into assertions vs. value-revealing prints.
@@
Conditioning on tasks that contain agent-written test artifact, we quantify how many feedback statements are encoded in those artifacts. We distinguish two signal types: (i) verification signals ($A$), i.e., assert statements that specify explicit checks, and (ii) observational signals ($P$), i.e., value-revealing print statements that expose runtime values or computed expressions. To ensure $P$ (prints) reflects observational feedback, we exclude pure-literal prints that emit only fixed strings (e.g., print("here")) and count only prints that expose runtime values, expressions, or execution results (e.g., print(obj.attr)). For each task $t$ with test artifacts $\mathcal{A}_{t}$, and for signal type $S\in\{A,P\}$, let $n^{S}_{t,a}$ be the number of signal statements of type $S$ in artifact $a\in\mathcal{A}_{t}$. We define the task-level signal totals:
@@
Resolved Unresolved Model #Tasks w/ tests Assertions per task ($\bar{N}^{A}$) Prints per task ($\bar{N}^{P}$) Total signals per task ($\bar{N}^{\mathrm{total}}$) #Tasks w/ tests Assertions per task ($\bar{N}^{A}$) Prints per task ($\bar{N}^{P}$) Total signals per task ($\bar{N}^{\mathrm{total}}$) claude-opus-4-5 314 5.16 25.00 30.16 101 5.36 25.61 30.97 gemini-3-pro-preview 235 1.45 4.34 5.79 73 1.62 5.04 6.66 kimi-k2-thinking 309 2.86 20.72 23.57 178 3.51 24.03 27.54 minimax-m2 302 7.37 34.06 41.43 191 4.66 43.09 47.76 deepseek-v3.2-reasoner 277 3.51 16.43 19.94 169 3.31 20.95 24.27
@@
Figure 2. Composition of feedback signals in agent-written tests across models. Value-revealing prints dominate over assertions for all models.
@@ Results.
Table 4 shows that, when agent-written tests are present, they can contain a substantial number of feedback statements per task (e.g., 19.94–47.76 total signals for several models). As shown in Figure 2, feedback is predominantly observational: for every model, value-revealing prints exceed assertions in the macro-average counts per task. Models differ markedly in overall signal volume. For example, minimax-m2 encodes the largest total signal counts (41.43–47.76), whereas gemini-3-pro-preview encodes far fewer (5.79–6.66), with other models in between. Across models, unresolved tasks tend to show slightly higher total signal counts, driven mainly by more value-revealing prints (e.g., deepseek-v3.2-reasoner: +4.33 total; minimax-m2: +6.33 total). Assertion counts are comparatively stable; notably, minimax-m2 has fewer assertions but more prints on unresolved tasks.
RQ2.1 Test Signal Amount: Key Takeaway When agents write tests, those tests often emit many runtime feedback statements per task, but the feedback is mostly observational: value-revealing prints consistently outnumber assertions across all models. Signal volume varies widely by model (e.g., minimax-m2 high vs. gemini-3-pro-preview low), and unresolved tasks show slightly more total signals mainly due to additional prints, while assertion counts remain relatively stable.
@@
To identify and categorize assertions, we implement a rule-based classifier over Python ASTs and map each extracted assertion to exactly one category. The classifier covers both native assert statements (e.g., assert a == b) and framework-provided assertion calls (e.g., self.assertEqual(a, b) in unittest). Concretely, for each test artifact, we parse the code into an AST and extract assertion events from: (i) native assert <expr> statements, and (ii) calls to framework assertion APIs. Some assert statements contain multiple checks in one line, combined with boolean operators (e.g., assert a > 0 and b == 1). In this example, a > 0 is a constraint check (C3) and b == 1 is an exact check (C4). For such compound assert statements, we decompose the expression into atomic checks and assign a single category by taking the highest category present under the ordering from C1 to C4, because the most specific check in the statement best reflects what the assertion is trying to enforce. Thus, assert a > 0 and b == 1 is labeled as C4.
@@
Model #Assertions C1 Sanity C2 Property C3 Relational C4 Exact # % # % # % # % claude-opus-4-5 2160 351 16.25% 807 37.36% 93 4.31% 909 42.08% gemini-3-pro-preview 458 76 16.59% 154 33.62% 36 7.86% 192 41.92% kimi-k2-thinking 1508 225 14.92% 622 41.25% 45 2.98% 616 40.85% minimax-m2 3117 618 19.83% 1291 41.42% 132 4.23% 1076 34.52% deepseek-v3.2-reasoner 1531 285 18.62% 537 35.08% 52 3.40% 657 42.91%
@@
Table 5 shows that models have broadly similar assertion-category distributions. Across all five models, most assertions fall into C2 Property and C4 Exact, while C3 Relational remains consistently uncommon. For four models (claude-opus-4-5, gemini-3-pro-preview, kimi-k2-thinking, and deepseek-v3.2-reasoner), C4 Exact accounts for roughly 41–43% of assertions (40.85–42.91%), and C2 Property accounts for roughly 34–41% (33.62–41.25%). minimax-m2 follows the same overall shape but allocates a smaller share to C4 Exact (34.52%) and larger shares to C1 Sanity (19.83%) and C2 Property (41.42%). Across all models, C3 Relational is rare (2.98–7.86%), with the highest proportion in gemini-3-pro-preview (7.86%). Overall, the consistent scarcity of C3 may have several practical explanations. First, relational or approximate checks can be more delicate to specify and maintain than local property checks or direct equality checks. Second, such checks may be less common in everyday unit-test patterns that models imitate during code generation. Third, for many SWE-bench issues, agents may find it more straightforward to write either local property checks (C2) or exact expected outputs (C4) once they have a candidate fix. We treat these distributions as descriptive of which assertion forms appear in agent-written tests, not as evidence of correctness or impact on task resolution.
RQ2.2 Assertion Categorization: Key Takeaway Across models, assertions in agent-written tests are dominated by two forms: checks of local properties (e.g., checking that a required attribute is present) and checks against exact expected values. Assertions that express relational or range-style constraints are consistently uncommon (only a small single-digit fraction of assertions across models). minimax-m2 has a lower share of exact-value checks and a higher share of sanity-guard and local-property checks than the other models, but the overall pattern is similar across models.
4.3. Summary of RQ2
RQ2 shifts from when agents test (RQ1) to what feedback their agent-written tests produce at execution time. First, when tests are present, they often emit many feedback statements per task, but the feedback is dominated by value-revealing prints rather than assertions: across all models, print signals consistently outnumber assert checks, and total signal volume varies widely by model (RQ2.1). Second, when agents do write assertions, models show broadly similar mixes of assertion forms: most assertions either check local properties (e.g., the presence or validity of an attribute or field) or check exact expected values, while relational or range-style constraints are rare (RQ2.2). Taken together, RQ2 characterizes agent-written tests primarily as an observational feedback channel: most test feedback comes from value-revealing prints. This naturally raises the next question: do these agent-written tests meaningfully affect task resolution? We address this in RQ3.
5. RQ3: Do Agent-Written Tests Truly Affect Task Resolution?
@@
In RQ1, we find a weak alignment between agent-written tests and the final task success in this high-autonomy setting. For example, gpt-5.2 almost never writes new test artifacts (3/500 tasks, 0.6%), yet it still resolves 71.8% of tasks. In contrast, claude-opus-4.5 writes at least one new test artifact in about 83% of tasks, but its resolution rate is only 2.6 percentage points higher (74.4%). RQ2 further shows that, when tests are written, most feedback comes from value-revealing prints rather than assert-based checks. These findings raise a direct question: Do writing tests truly affect task resolution outcomes, and at what cost?
@@
  • RQ3.1 (Outcome impact): If we encourage or discourage agents to write tests, how do task resolution outcomes change?
  • RQ3.2 (Efficiency impact): If we encourage or discourage agents to write tests, how do API calls and token usage change?
@@
To isolate the effect of agent-written tests, we design two complementary intervention experiments: (i) encouraging agents to write tests, and (ii) discouraging agents from writing new test files. We choose models for each setup based on their baseline test-writing rate observed in RQ1 (Table 1), defined as the fraction of tasks where the agent writes test artifacts.
For the encourage test writing setup, we focus on low test-writing models and medium test-writing models, so there is meaningful headroom for increasing test creation. Specifically, we include gpt-5.2 (0.6%), an extreme low test-writing model in RQ1 with near-zero test creation. We also include gemini-3-pro-preview (61.1%), a medium test-writing model whose baseline test creation is already substantial but still leaves room for further increase.
For the discourage test writing setup, we start from high test-writing models that write tests in the vast majority of tasks in RQ1: four models show consistently high test-writing rates (83.0%–98.6%; Table 1). Due to budget constraints, we select two representatives from this group: kimi-k2-thinking (97.4%) and deepseek-v3.2-reasoner (89.2%).
Concretely, we use two prompt variants:
  • Encourage writing tests: for gpt-5.2 and gemini-3-pro-preview, we append an instruction to write at least one runnable new test file (a file whose name starts with test_ or ends with _test.py), separate from the repository’s existing tests.
  • Discourage writing tests: for kimi-k2-thinking and deepseek-v3.2-reasoner, we (i) remove the sentence “Test edge cases to ensure your fix is robust” and (ii) append an instruction to not write any new test files or scripts; robustness and edge cases should be handled using reasoning and code inspection only.
By comparing the performance of each agent under these revised prompts against their original performance, we isolate the specific impact of test writing. Consequently, the observed changes in behaviors or outcomes can be directly attributed to the agent-written tests.
5.1. RQ3.1 Does encouraging or discouraging test writing change task resolution?
@@
To answer whether encouraging or discouraging test writing changes task resolution, for each model, we compare each task under two conditions: the baseline run (under standard mini-SWE-agent prompt) and the intervention run (under our revised prompt). Specifically, we record two features: whether the run creates at least one new test artifact (No test vs. Has test) and whether the patch successfully resolve the issue (Fail vs. Success). Then, we analyze how these two features change from the baseline run to the intervention run. This respectively results in four possible transition groups for both test writing (No $\rightarrowNo test$, No $\rightarrowHas test$, Has $\rightarrowNo test$, Has $\rightarrowHas test$) and task outcome (Fail $\rightarrow$ Success, Success $\rightarrow$ Fail, Stable Success, and Stable Fail). To visualize the relationship between these shifts, we represent the results using a transition matrix. In this matrix, the rows represent the change in test-writing behavior, while the columns represent the change in task outcomes. This structure allows us to pinpoint the exact impact of our intervention. For instance, the intersection of No test $\rightarrow$ Has test and Fail $\rightarrow$ Success represents the instances where encouraging test-writing directly led to a breakthrough in resolving the task.
@@ Results.
Our prompt interventions substantially change whether models write test artifacts, but these shifts rarely translate into outcome changes. As shown in Table 6, the encourage test writing prompt flips test status for the low test-writing model gpt-5.2 and the medium test-writing model gemini-3-pro-preview: 64.4% and 37.0% of tasks transition from No test to Has test, respectively. Conversely, the discourage test writing prompt removes tests at scale for the high test-writing models kimi-k2-thinking and deepseek-v3.2-reasoner, moving 68.4% and 75.2% of tasks from Has test to No test.
Despite these large test-status shifts, resolution outcomes are largely stable. Figure 3 shows that, across models, an average of 83.2% of tasks keep the same final resolution result after the intervention. Table 6 further indicates that even when test status flips, success rates change only slightly. For example, for deepseek-v3.2-reasoner, discouraging test writing removes tests in 376 tasks but yields a net decrease of only 20 resolved tasks, a small change relative to the behavioral shift. Overall, changing how often a model writes test artifacts appears to be a weak lever for shifting task outcomes in this setting.
RQ3.1: Outcomes vs. Test-Status Changes Discouraging test writing removes agent-written tests at scale (at least 68% of tasks for kimi-k2-thinking and deepseek-v3.2-reasoner). Yet these large behavioral shifts rarely translate into outcome shifts: across models, 83.2% of tasks keep the same final resolution result after the intervention. Even when tests are induced, success does not increase correspondingly, suggesting that simply writing more test artifacts provides little leverage for fixing the underlying issue. Conversely, suppressing tests for hundreds of tasks does not cause outcomes to collapse; for deepseek-v3.2-reasoner, removing tests in 376 tasks yields only 20 fewer successes.
@@
We further analyze the following three metrics based on the trajectories generated in RQ3.1: (i) average API calls per task, (ii) average input tokens per task, and (iii) average output tokens per task.
Table 7. API calls and token usage under baseline vs. encourage-tests / discourage-tests conditions.
Model Condition Tasks resolved Avg API Calls Avg Input Tokens Avg Output Tokens gpt-5.2 Baseline 359 (71.8%) 19.76 242,855 24,550 Encourage writing tests 359 (71.8%) 20.84 264,762 29,415 Change +0 (+0.0%) +1.08 (+5.5%) +21,907 (+9.0%) +4,866 (+19.8%) gemini-3-pro-preview Baseline 371 (74.2%) 40.33 666,096 11,114 Encourage writing tests 366 (73.2%) 39.21 641,307 10,943 Change -5 (-1.0%) -1.11 (-2.8%) -24,789 (-3.7%) -171 (-1.5%) kimi-k2-thinking Baseline 317 (63.4%) 46.82 668,449 14,895 Discourage writing tests 304 (60.8%) 30.25 340,689 8,468 Change -13 (-2.6%) -16.57 (-35.4%) -327,760 (-49.0%) -6,427 (-43.1%) deepseek-v3.2-reasoner Baseline 300 (60.0%) 46.40 637,297 52,120 Discourage writing tests 291 (58.2%) 35.06 427,780 44,823 Change -9 (-1.8%) -11.35 (-24.5%) -209,518 (-32.9%) -7,297 (-14.0%)
Note. Changes are computed as (Condition - Baseline). Encourage writing tests is applied to gpt-5.2 and gemini-3-pro-preview; Discourage writing tests is applied to kimi-k2-thinking and deepseek-v3.2-reasoner.
@@
Table 7 reports task outcomes (Tasks resolved) and efficiency metrics (API calls and tokens) under the baseline and the two intervention conditions. Overall, the interventions have only marginal impact on resolution rates, but they can noticeably reshape efficiency—in ways that depend on each model’s baseline test-writing propensity.
For the encourage test writing setup, the effects differ between the low test-writing model gpt-5.2 and the medium test-writing model gemini-3-pro-preview. For gpt-5.2, encouraging test writing increases overhead: API calls rise by 5.5% and output tokens by 19.8%, while the resolution rate remains unchanged from baseline. In the medium test-writing model gemini-3-pro-preview, we observe very small shifts in overall usage (e.g., $1.5\%$ completion tokens). A likely reason is that this model already writes tests in about 61.6% of baseline tasks, leaving limited headroom for the prompt to flip tasks from No test to Has test; accordingly, the net change in testing-related interaction is small on average. Moreover, because the model is nondeterministic, even when a task contains tests in both baseline and encouraged runs, the amount of testing (and related interaction) can differ slightly between the paired runs; these small within-task fluctuations average out, keeping aggregate cost differences modest and making a small decrease reasonable.
The most striking efficiency shifts appear in the discourage test writing setup for the high test-writing models kimi-k2-thinking and deepseek-v3.2-reasoner. Discouraging test writing yields large reductions in resource consumption, with input tokens dropping by 49.0% and 32.9%, respectively. For kimi-k2-thinking, average API calls fall by 35.4%, cutting interaction workload by over a third. Notably, these efficiency gains come with only small decreases in success rates (2.6% for kimi-k2-thinking and 1.8% for deepseek-v3.2-reasoner), suggesting that a substantial portion of the baseline interaction budget spent on test writing for these models has limited marginal benefit for task resolution in this setting.
RQ3.2: Cost changes are larger than outcome changes Changing whether agents write tests has little effect on task resolution, but it can noticeably change efficiency. More tests increases overhead for low test-writing model models without improving outcomes: for gpt-5.2, API calls rise by 5.5% and output tokens by 19.8% while the resolution rate stays unchanged. In contrast, suppressing tests yields large efficiency gains for high test-writing models: input tokens drop by 49.0% (kimi-k2-thinking) and 32.9% (deepseek-v3.2-reasoner), and kimi-k2-thinking reduces API calls by 35.4%. These savings come with only small success drops (2.6% and 1.8%), implying that most test-writing overhead in these models delivers low marginal utility for producing a successful final patch.

### v2

@@ 4. RQ2: What Feedback Signals Do Agent-Written Tests Provide?
RQ2 conditions on tasks that write at least one test artifact and reports descriptive summaries of three aspects of test feedback:
  • Signal counts (RQ2.1): assertion vs. value-revealing print counts in agent-written tests.
@@
  • Print types (RQ2.3): what value-revealing prints typically inspect, using a three-type categorization.
RQ2.1 stays at the task level and measures how much feedback agent-written tests emit overall. RQ2.2 and RQ2.3 then move to the statement level, categorizing assertion and value-revealing print statements separately.
@@
Conditioning on tasks that contain agent-written test artifact, we quantify how many feedback statements are encoded in those artifacts. We distinguish two signal types: (i) verification signals (AA), i.e., assert statements that specify explicit checks, and (ii) observational signals ($P$), i.e., value-revealing print statements that expose runtime values or computed expressions. To ensure $P$ (prints) reflects observational feedback, we exclude pure-literal prints that emit only fixed strings (e.g., print("here")) and count only prints that expose runtime values, expressions, or execution results (e.g., print(obj.attr)). For each task $t$ with test artifacts $\mathcal{A}_{t}$, and for signal type $S\in\{A,P\}$, let $n^{S}_{t,a}$ be the number of signal statements of type $S$ in artifact $a\in\mathcal{A}_{t}$. We define the task-level signal totals:
@@
Resolved Unresolved Model #Tasks w/ tests Assertions per task ($\bar{N}^{A}$) Prints per task ($\bar{N}^{P}$) Total signals per task ($\bar{N}^{\mathrm{total}}$) #Tasks w/ tests Assertions per task ($\bar{N}^{A}$) Prints per task ($\bar{N}^{P}$) Total signals per task ($\bar{N}^{\mathrm{total}}$) Claude 4.5 314 5.16 25.00 30.16 101 5.36 25.61 30.97 Gemini 3 Pro 235 1.45 4.34 5.79 73 1.62 5.04 6.66 Kimi K2-T 309 2.86 20.72 23.57 178 3.51 24.03 27.54 MiniMax M2 302 7.37 34.06 41.43 191 4.66 43.09 47.76 DeepSeek v3.2-R 277 3.51 16.43 19.94 169 3.31 20.95 24.27
@@
Figure 2. Composition of feedback signals in agent-written tests across models.
@@ Results.
Table 4 shows that, when agent-written tests are present, they can contain a substantial number of feedback statements per task. As shown in Figure 2, feedback is predominantly observational: for every model, value-revealing prints exceed assertions in the macro-average counts per task. Signal volume also varies markedly by model, from the lower totals of Gemini 3 Pro to the much higher totals of MiniMax M2. Across models, unresolved tasks tend to show slightly higher total signal counts, driven mainly by more value-revealing prints, while assertion counts remain comparatively stable.
RQ2.1 Test Signal Amount: Key Takeaway When agents write tests, the feedback is mostly observational: value-revealing prints consistently outnumber assertions, although total signal volume still varies markedly by model.
@@
To identify and categorize assertions, we implement a rule-based classifier over Python ASTs and map each assertion to exactly one category. The classifier covers both native assert statements (e.g., assert a == b) and framework-provided assertion calls (e.g., self.assertEqual(a, b) in unittest). Concretely, for each test artifact, we parse the code into an AST and extract assertion events from: (i) native assert <expr> statements, and (ii) calls to framework assertion APIs. Some assert statements contain multiple checks in one line with boolean operators (e.g., assert a > 0 and b == 1). In this example, a > 0 is a constraint check (C3) and b == 1 is an exact check (C4). For such compound assert statements, we decompose the expression into atomic checks and assign a single category by taking the highest category under the ordering from C1 to C4, because the most specific check in the statement best reflects what the assertion is trying to enforce. Thus, assert a > 0 and b == 1 is labeled C4.
@@
Model #Assertions C1 Sanity C2 Property C3 Relational C4 Exact # % # % # % # % Claude 4.5 2160 351 16.25% 807 37.36% 93 4.31% 909 42.08% Gemini 3 Pro 458 76 16.59% 154 33.62% 36 7.86% 192 41.92% Kimi K2-T 1508 225 14.92% 622 41.25% 45 2.98% 616 40.85% MiniMax M2 3117 618 19.83% 1291 41.42% 132 4.23% 1076 34.52% DeepSeek v3.2-R 1531 285 18.62% 537 35.08% 52 3.40% 657 42.91%
@@
Table 5 shows that models have broadly similar assertion-category distributions. Across all five models, most assertions fall into C2 Property and C4 Exact, while C3 Relational remains consistently uncommon. For four models (Claude Opus 4.5, Gemini 3 Pro, Kimi K2 Thinking, and DeepSeek v3.2 Reasoner), C4 Exact accounts for roughly 41–43% of assertions (40.85–42.91%), and C2 Property accounts for roughly 34–41% (33.62–41.25%). MiniMax M2 follows the same overall shape but allocates a smaller share to C4 Exact (34.52%) and larger shares to C1 Sanity (19.83%) and C2 Property (41.42%). Across all models, C3 Relational is rare (2.98–7.86%), with the highest proportion in Gemini 3 Pro (7.86%). This scarcity suggests that agents more often fall back on local property checks (C2) or exact expected outputs (C4), while relational or approximate constraints may be harder to specify and less common in the test patterns models imitate. We treat these distributions as descriptive of which assertion forms appear in agent-written tests, not as evidence of correctness or impact on task resolution.
RQ2.2 Assertion Categorization: Key Takeaway Across models, assertions are dominated by property checks and exact-value checks, whereas relational or range-style constraints remain uncommon.
4.3. RQ2.3 Print categorization: What do value-revealing prints typically inspect?
Goal and measurements.
Since value-revealing prints substantially outnumber assertions, we further inspect what these prints expose in practice. This helps clarify whether agent-written tests are primarily used to inspect values, inspect structural summaries, or surface execution status, thereby refining our interpretation of tests as observational debugging tools. We group value-revealing prints into three categories:
  • P1 Value / content inspection. The print inspects a concrete runtime value or content, such as a returned output, an intermediate result, an object field, or part of a generated string. This category is used when the print is meant to reveal what the program produced, rather than a structural summary or an error/status signal; e.g., print(add(1, 2)).
  • P2 Structural summary inspection. The print reports an aggregate structural summary, such as length, size, shape, number of items, or emptiness. This category is reserved for prints that summarize how much data is present or what structure it has, rather than showing the content itself; e.g., print(len(items)).
  • P3 Exception / execution-status signals. The print surfaces an exception, an error message, or a coarse execution-status indicator, such as a success/failure flag. This category is used when the print primarily indicates what happened during execution, rather than the program’s computed content; e.g., except Exception as exc: print(exc).
We implement a rule-based classifier over Python ASTs. For each parseable test artifact, we extract value-revealing print(...) calls, exclude pure-literal or formatting-only prints, and assign each print to one of the three categories above using deterministic category rules.
Table 6. Value-revealing print category distribution by model.
Model #Prints P1 Value/Content P2 Structural P3 Exception/Status # % # % # % Claude 4.5 7919 6136 77.48% 274 3.46% 1509 19.06% Gemini 3 Pro 1352 942 69.67% 72 5.33% 338 25.00% Kimi K2-T 8909 6556 73.59% 584 6.56% 1769 19.86% MiniMax M2 15685 11003 70.15% 921 5.87% 3761 23.98% DeepSeek v3.2-R 7495 5709 76.17% 280 3.74% 1506 20.09%
Results.
Table 6 shows a stable pattern across models. P1 Value / content inspection dominates for every model (69.67–77.48%), indicating that most prints are used to inspect concrete outputs, intermediate values, or object contents. P3 Exception / execution-status signals form the second-largest category (19.06–25.00%), whereas P2 structural-summary inspection remains comparatively uncommon (3.46–6.56%). Overall, value-revealing prints are used mainly to inspect runtime values and coarse execution outcomes rather than encode strong pass/fail criteria, reinforcing our interpretation of agent-written tests as observational debugging tools.
RQ2.3 Print Categorization: Key Takeaway Across models, most value-revealing prints are used for value/content inspection, with exception/execution-status signals as a distant second. Structural summaries are much less common, reinforcing that these prints function primarily as observational debugging probes rather than strong correctness oracles.
$4.$ Summary of RQ2
RQ2 shows that agent-written tests function primarily as an observational feedback channel: value-revealing prints dominate, and the assertions that do appear are concentrated in local-property and exact-value checks. This naturally raises the next question: do these agent-written tests meaningfully affect task resolution?
5. RQ3: How Does Prompting Test Writing Change Observed Outcomes and Costs?
@@
In RQ1, we find a weak alignment between agent-written tests and the final task success in this high-autonomy setting. For example, GPT-5.2 almost never writes new test artifacts (3/500 tasks, 0.6%), yet it still resolves 71.8% of tasks. In contrast, Claude Opus 4.5 writes at least one new test artifact in about 83% of tasks, but its resolution rate is only 2.6 percentage points higher (74.4%). RQ2 further shows that, when tests are written, most feedback comes from value-revealing prints rather than assert-based checks. These findings raise a direct question: when prompts change whether agents write tests, how do the observed task outcomes and costs change in this setting?
@@
  • RQ3.1 (Outcome changes): If we encourage or discourage agents to write tests, how do observed task resolution outcomes change?
  • RQ3.2 (Efficiency changes): If we encourage or discourage agents to write tests, how do API calls and token usage change?
@@
To probe how prompting test writing changes observed behavior under the same scaffold, we design two complementary intervention experiments: (i) encouraging agents to write tests, and (ii) discouraging agents from writing new test files. We choose models for each setup based on their baseline test-writing rate observed in RQ1 (Table 1), defined as the fraction of tasks where the agent writes test artifacts.
For the encourage test writing setup, we focus on low test-writing models and medium test-writing models, so there is meaningful headroom for increasing test creation. We do not include already test-heavy models in this setup, because their baseline test-writing rates leave little room for a meaningful upward shift. Specifically, we include GPT-5.2 (0.6%), an extreme low test-writing model in RQ1 with near-zero test creation. We also include Gemini 3 Pro (61.1%), a medium test-writing model whose baseline test creation is already substantial but still leaves room for further increase.
For the discourage test writing setup, we start from high test-writing models that write tests in the vast majority of tasks in RQ1: four models show consistently high test-writing rates (83.0%–98.6%; Table 1). Due to budget constraints, we select two representatives from this group: Kimi K2 Thinking (97.4%) and DeepSeek v3.2 Reasoner (89.2%).
Concretely, we start from the original mini-SWE-agent prompt and make small, targeted prompt edits to create two variants:
  • Encourage writing tests: for GPT-5.2 and Gemini 3 Pro, we append prompt instructions to write at least one runnable new test file (a file whose name starts with test_ or ends with _test.py), separate from the repository’s existing tests.
  • Discourage writing tests: for Kimi K2 Thinking and DeepSeek v3.2 Reasoner, we remove the default testing-related prompt cue from mini-SWE-agent and append prompt instructions not to write any new test files or scripts.
The exact texts of the original prompt and both revised variants are included in the experiment_prompts folder. [^2] Specifically, the files are mini_swe_agent_original_prompt.yaml, encourage_write_tests_prompt.yaml, and discourage_write_tests_prompt.yaml. We compare each revised prompt against its baseline to measure how changing the prompt is accompanied by changes in test writing, outcomes, and costs.
5.1. RQ3.1 How does encouraging or discouraging test writing change observed task outcomes?
@@
[…]
@@ Results.
Our prompt interventions substantially change whether models write test artifacts, but these shifts rarely translate into outcome changes. As shown in Table 7, the encourage test writing prompt flips test status for the low test-writing model GPT-5.2 and the medium test-writing model Gemini 3 Pro: 64.4% and 37.0% of tasks transition from No test to Has test, respectively. Conversely, the discourage test writing prompt removes tests at scale for the high test-writing models Kimi K2 Thinking and DeepSeek v3.2 Reasoner, moving 68.4% and 75.2% of tasks from Has test to No test.
Despite these large test-status shifts, resolution outcomes are largely stable. Figure 3 shows that, across models, an average of 83.2% of tasks keep the same final resolution result after the intervention. Table 7 further indicates that even when test status flips, success rates change only slightly. The exact McNemar tests reinforce this descriptive picture: none of the four models shows a statistically significant baseline-vs.-intervention outcome shift (all $p>0.05$). For example, for DeepSeek v3.2 Reasoner, discouraging test writing removes tests in 376 tasks but yields a net decrease of only 20 resolved tasks, a small change relative to the behavioral shift. Overall, changing how often a model writes test artifacts appears to be a weak lever for shifting task outcomes in this setting.
5.1.1. Exploring Issue Types for Which Tests May Help
[…]
Results.
We find 8 issues that appear in both categories: astropy-13236; django-11790, django-13401, django-13512, and django-14493; pylint-7080; scikit-learn-14629; and sympy-21612. We manually read the problem statements of these 8 issues and used GPT-5.4 to help summarize recurring themes. This exploratory pass suggests three recurring properties: they are mostly small-to-moderate correctness bugs, often involve explicit reproduction conditions or edge-case triggers, and frequently require checking precise expected behavior or semantics.
These observations are necessarily limited. SWE-bench Verified contains 500 tasks in total, the benchmark does not provide an official issue taxonomy, and our cross-category overlap contains only 8 instances. We therefore present this subsection as an exploratory step beyond the main intervention result, intended to surface plausible test-sensitive candidates rather than to establish definitive issue categories. A broader taxonomy-based analysis is left to future work.
RQ3.1: Outcomes vs. Test-Status Changes Prompt interventions can flip test-writing status at scale, yet most tasks keep the same final outcome. Even large induced or suppressed shifts in test writing therefore appear to be a weak lever on task resolution in this setting.
@@
We further analyze the following three metrics based on the trajectories generated in RQ3.1: (i) average API calls per task, (ii) average input tokens per task, and (iii) average output tokens per task. For each model and metric, we compare the intervention run against the baseline run on the same 500 tasks. To quantify whether the observed efficiency changes are directionally robust, we additionally compute two-sided Wilcoxon signed-rank tests and 95% paired bootstrap confidence intervals for the task-level difference (Condition - Baseline).
Table 8. API calls and token usage under baseline vs. test-encouragement / test-discouragement conditions.
Model Condition Tasks resolved Avg API Calls Avg Input Tokens Avg Output Tokens GPT-5.2 Baseline 359 (71.8%) 19.76 242,855 24,550 Encourage tests 359 (71.8%) 20.84 264,762 29,415 Change -+0 (+0.0%) -+1.08 (+5.5%)∗∗∗ -+21,907 (+9.0%)∗∗∗ -+4,866 (+19.8%)∗∗∗ Gemini 3 Pro Baseline 371 (74.2%) 40.33 666,096 11,114 Encourage tests 366 (73.2%) 39.21 641,307 10,943 Change -5 (-1.0%) -1.11 (-2.8%) -24,789 (-3.7%) -171 (-1.5%) Kimi K2-T Baseline 317 (63.4%) 46.82 668,449 14,895 Discourage tests 304 (60.8%) 30.25 340,689 8,468 Change -13 (-2.6%) -16.57 (-35.4%)∗∗∗ -327,760 (-49.0%)∗∗∗ -6,427 (-43.1%)∗∗∗ DeepSeek v3.2-R Baseline 300 (60.0%) 46.40 637,297 52,120 Discourage tests 291 (58.2%) 35.06 427,780 44,823 Change -9 (-1.8%) -11.35 (-24.5%)∗∗∗ -209,518 (-32.9%)∗∗∗ -7,297 (-14.0%)∗∗∗
Note. Changes are computed as (Condition - Baseline). Encourage tests is applied to GPT-5.2 and Gemini 3 Pro; Discourage tests is applied to Kimi K2-T and DeepSeek v3.2-R. $^{***}p<0.001$ by paired Wilcoxon signed-rank test; paired bootstrap CIs are summarized in the text.
@@
Table 8 shows that the interventions have only marginal impact on resolution rates, but can noticeably reshape efficiency. Under encourage test writing, the low test-writing model GPT-5.2 incurs higher overhead (+5.5% API calls; +19.8% output tokens) without any gain in resolution. The paired analyses support this increase across all three metrics: API calls rise by +1.08 per task (95% CI [0.39, 1.75], Wilcoxon $p<0.001$), input tokens by +21,907 ([3,954, 39,598], $p<0.001$), and output tokens by +4,866 ([3,041, 6,704], $p<0.001$). In contrast, the medium test-writing model Gemini 3 Pro changes little overall, and none of its paired differences reaches conventional significance across API calls, input tokens, or output tokens (all 95% CIs cross zero). The most striking shifts appear under discourage test writing for the high test-writing models Kimi K2 Thinking and DeepSeek v3.2 Reasoner: input tokens drop by 49.0% and 32.9%, respectively, and Kimi K2 Thinking also reduces API calls by 35.4%. These reductions are directionally robust in the paired analyses: for both models, all three mean differences are negative, all 95% bootstrap CIs exclude zero, and all Wilcoxon tests give $p<0.001$. These paired results therefore strengthen the interpretation that suppressing test writing yields substantial efficiency savings in these two high test-writing models, whereas encouraging test writing produces model-dependent and much less stable cost effects.
RQ3.2: Cost changes are larger than outcome changes Changing whether agents write tests has much larger effects on efficiency than on task resolution. Paired Wilcoxon tests and bootstrap confidence intervals support the large discourage-setting savings, while encourage-setting effects are robust for GPT-5.2 but small and statistically unclear for Gemini 3 Pro.